# Benchmarking LLMs in Web API Integration Tasks

### v2

@@ 3.1. Results
Mean arg. val. cond. acc. (e) 000.53 000.68 000.79 000.72 000.76 000.00 000.57 000.84 000.00 000.95
Table 4. Evaluation results for argument completion.
@@
Mean arg. val. cond. acc. (e) 000.71 000.83 000.83 000.83 000.86 000.78 000.65 000.89 000.67 000.95
3.2. Discussion
The models perform comparatively well in predicting the HTTP method, which is expected since there is only a small set of available methods, with GET and POST being by far the most common ones. In contrast, URLs and arguments are challenging to predict due to their unrestricted nature. The high argument precision can be attributed to the Authorization header argument, which is required my most endpoints and therefore easy to predict. The relatively high argument value conditional accuracy indicates that identifying the correct argument is harder for the models than assigning the correct value to it once it is found.
An analysis of the generated implementations shows that those models that got nearly zero implementations correct in the full completion setup failed to understand the task given to them and refused to properly continue the starter code, leading to incomplete and thus non-executable code. It may be possible to solve this issue through prompt engineering, but we preferred to use a single prompt consistently for all models.
An additional finding is that larger models are not always better. In multiple model families (DeepSeek-Coder, Qwen2.5-Coder, Code Llama), we observe that the medium-sized variant performs worse than both the smaller and larger variants. Lastly, breaking down the evaluation results by API (cf. tables in the appendix), we see that each model is particularly good or bad at different APIs, which likely results from the prevalence of the respective APIs in the model’s training data.
3.3. Limitations and Threads to Validity
1) Real-world API integration code is diverse and complex. To be able to evaluate API invocations automatically, we had to apply some restrictions and simplifications when creating the tasks in our dataset, as described in Section 2.1. For instance, there is limited program context before the API invocation. Additionally, we evaluate only the correctness of the outgoing request, not how the incoming response is handled. These limitations do, however, not interfere with our investigation of LLM’s ability to recall information about correct API usage—if this information is memorized, they should be able to recall it in our evaluation and real-world settings.
@@ 4. Related Work
Since APIs are such a universal concept, it is important to distinguish between different settings in which APIs are being used and understand their different characteristics when discussing related work. We divide APIs roughly into the following categories: (1) general web APIs(this work), (2) domain-specific web APIs, (3) SDK-wrapped web APIs, (4) local function APIs, and (5) tool APIs for AI agents.
@@
2) Su et al. (Su et al., 2017, 2018) and Hosseini et al. (Hosseini et al., 2021) investigated natural language interfaces to domain-specific web APIs following the Open Data Protocol [^12] https://docs.oasis-open.org/odata/odata/v4.0/odata-v4.0-part1-protocol.html (OData). Their work has a narrow focus on data-centric applications (mostly within the Microsoft ecosystem) and an SQL-like syntax. In contrast, our benchmark targets OpenAPI, a very general and flexible API standard, making the task more open-ended and challenging. Also, instead of ad-hoc API queries by a user, we strive for production-ready API integration code implemented in JavaScript.
3) Works like Dialog2API (Shu et al., 2022) and CloudAPIBench (Jain et al., 2024) focused on SDK-wrapped web API invocations (e.g., for AWS), treating web APIs like local libraries. Both approaches do not generalize due to SDK heterogeneity. In contrast, our benchmark relies on a unified RESTful interface for any web API. This eliminates SDK dependency, enabling direct assessment of models’ core API comprehension rather than SDK-specific knowledge. By standardizing interactions across diverse APIs, we remove evaluation biases and isolate model capabilities from implementation quirks of individual SDKs.
4) Local function APIs are the most common kind of API in typical codebases and studied in many works, e.g., (Zan et al., 2023; Eghbali and Pradel, 2024; Dutta et al., 2024). Invoking a function requires specifying the function name and one argument list. In contrast, general web API invocations are subject to more complex syntax and semantics than such local function calls, as already mentioned in Section 1 (Scope).
5) There is a large body of work concerned with enabling LLM-based AI agents to use tools to interact with their environment (Yehudai et al., 2025). These tools range from simple utilities like calculators (Schick et al., 2023) to web services that enable AI agents to perform tasks such as weather queries or restaurant booking (Song et al., 2023; Li et al., 2023b; Qin et al., 2024). At first sight, this body of work seems related to research on automatic generation of integration code, as tools typically expose a function-like interface. However, both input and output modalities differ between these settings, as well as their non-functional requirements, so benchmarks are not directly transferable to other kinds of API invocations.
@@
Since functional abstractions are essential in programming, virtually every code dataset includes (function) API invocations. Code datasets involving web APIs can be found in the fields of natural language interfaces to web APIs (Su et al., 2017; Hosseini et al., 2021), SDK-based API invocations (Shu et al., 2022; Jain et al., 2024), and tool-using AI agents (Patil et al., 2023; Song et al., 2023; Li et al., 2023b; Qin et al., 2024; Yan et al., 2024). We are not aware of datasets specifically about generating API integration code to invoke REST APIs like we do.
@@
The correctness of AI-generated code is traditionally measured either using exact matches or syntax-based similarity, which both correlate weakly with actual correctness (Evtikhiev et al., 2023), or using functional testing (Chen et al., 2021), which is not directly transferable to web API invocations, as discussed in Section 2.3. This is why we developed a novel approach described in Section 2. Evaluation methods in related works on API invocations mostly rely on variations of the aforementioned standard approaches. Some works use humans (Ding et al., 2025) or LLMs (Qin et al., 2024) as judges, and others execute the API invocation and verify the state of the environment (Shu et al., 2022; Yan et al., 2024). We are not aware of an evaluation method able to provide results as fine-grained as ours.
@@
LLMs display a remarkable capacity to memorize information from their training data—which can be beneficial when it comes to generating factual responses, but also a problem when it comes to leaking private information from the training data. Studies have shown that LLMs memorize, among other things, method definitions and method calls (Yang et al., 2024), but also that they struggle to suggest the correct APIs (Zhuo et al., 2023). These findings are in line with our observation that LLMs seem to know many endpoint URLs and argument names, but are often unable to assemble them to a specification-compliant API invocation that solves the task in question.
If LLMs are unable to recall memorized facts, they tend to hallucinate information that is not grounded in reality. This is a common error of LLMs, and studies have shown that it also exists in code generation, where, among other things, models may hallucinate function names and arguments (Liu et al., 2024; Dou et al., 2024; Tambon et al., 2025). Our study confirms that these findings also hold for web API endpoints and their arguments, where a high amount of hallucinations can be observed as well.
@@
Correct usage of APIs is challenging also for humans, as indicated by several studies on API usability issues (Robillard, 2009; Rauf et al., 2019) and real-world API misuses (Bae et al., 2014; Li et al., 2021; Wang and Zhao, 2023). Several (static analysis) methods have been proposed for detecting (and mitigating) API misuses, e.g., (Bae et al., 2014; Binder et al., 2022; Wang and Zhao, 2023). They are, however, primarily designed to aid humans, not LLMs. Using static analyses for evaluation purposes is generally conceivable, but they can only detect whether an API invocation is legal or illegal, not whether it also solves the given task.
5. Conclusion and Future Work
@@
In this paper, we examined the challenges faced by large language models when generating reliable web API invocation code. We introduced a novel dataset and evaluation pipeline to systematically assess LLM performance in this domain. Our experiments with open-source LLMs revealed significant limitations: Only 30% (full completion) and 40% (argument completion) of the produced samples were completely right. Endpoints and arguments are often hallucinated due to insufficient recall of knowledge on API usage.
@@
Overall, our work underscores the need for integrating quality assurance techniques into AI-driven code generation workflows to ensure correctness and reliability. Our benchmark lays the foundation for future work to investigate the impact of different approaches—such as retrieval-augmented generation (Lewis et al., 2020), fine-tuning, reasoning, feedback loops, and constrained decoding (Hokamp and Liu, 2017; Deutsch et al., 2019)—on the correctness and specification-compliance of generated web API invocation code. While our evaluation pipeline is specialized on JavaScript and Axios, our dataset and methodology is generic and can be transferred to other languages and libraries.
Acknowledgements.

### v3

@@ III-B Discussion
Mean argument value conditional accuracy (e) 000.53 000.68 000.79 000.72 000.76 000.00 000.57 000.84 000.00 000.95
TABLE IV: Evaluation Results for Argument Completion
@@
Mean argument value conditional accuracy (e) 000.71 000.83 000.83 000.83 000.86 000.78 000.65 000.89 000.67 000.95
III-C Limitations and Threats to Validity
1) Real-world API integration code is diverse and complex. To be able to evaluate API invocations automatically, we had to apply some restrictions and simplifications when creating the tasks in our dataset, as described in Section II-A. For instance, there is limited program context before the API invocation. Additionally, we evaluate only the correctness of the outgoing request, not how the incoming response is handled. These limitations do, however, not interfere with our investigation of LLM’s ability to recall information about correct API usage—if this information is memorized, they should be able to recall it in our evaluation as well as real-world settings.
@@ IV Related Work
Since APIs are such a universal concept, it is important to distinguish between different settings in which APIs are being used and understand their different characteristics when discussing related work. We divide APIs roughly into the following categories: 1. general web APIs(this work), 2. domain-specific web APIs, 3. SDK-wrapped web APIs, 4. local function APIs, and 5. tool APIs for AI agents.
@@
2) Su et al. [27, 28] and Hosseini et al. [29] investigated natural language interfaces to domain-specific web APIs following the Open Data Protocol [^12] https://docs.oasis-open.org/odata/odata/v4.0/odata-v4.0-part1-protocol.html (OData). Their work has a narrow focus on data-centric applications (mostly within the Microsoft ecosystem) and an SQL-like syntax. In contrast, our benchmark targets OpenAPI, a very general and flexible API standard, making the task more open-ended and challenging. Also, instead of ad-hoc API queries by a user, we strive for production-ready API integration code implemented in JavaScript.
3) Works like Dialog2API [30] and CloudAPIBench [31] focused on SDK-wrapped web API invocations (e.g., for AWS), treating web APIs like local libraries. Both approaches do not generalize due to SDK heterogeneity. In contrast, our benchmark relies on a unified RESTful interface for any web API. This eliminates SDK dependency, enabling direct assessment of models’ core API comprehension rather than SDK-specific knowledge. By standardizing interactions across diverse APIs, we remove evaluation biases and isolate model capabilities from implementation quirks of individual SDKs.
4) Local function APIs are the most common kind of API in typical codebases and studied in many works, e.g., [32, 33, 34]. Invoking a function requires specifying the function name and one argument list. In contrast, general web API invocations are subject to more complex syntax and semantics than such local function calls, as already mentioned in Section I (Scope).
5) There is a large body of work concerned with enabling LLM-based AI agents to use tools to interact with their environment [35]. These tools range from simple utilities like calculators [36] to web services that enable AI agents to perform tasks such as weather queries or restaurant booking [37, 38, 39]. At first sight, this body of work seems related to research on automatic generation of integration code, as tools typically expose a function-like interface. However, both input and output modalities differ between these settings, as well as their non-functional requirements, so benchmarks are not directly transferable to other kinds of API invocations.
@@
Since functional abstractions are essential in programming, virtually every code dataset includes (function) API invocations. Code datasets involving web APIs can be found in the fields of natural language interfaces to web APIs [27, 29], SDK-based API invocations [30, 31], and tool-using AI agents [40, 37, 38, 39, 41]. We are not aware of datasets specifically about generating API integration code to invoke REST APIs like we do.
@@
The correctness of AI-generated code is traditionally measured either using exact matches or syntax-based similarity, which both correlate weakly with actual correctness [23], or using functional testing [24], which is not directly transferable to web API invocations, as discussed in Section II-C. This is why we developed a novel approach described in Section II. Evaluation methods in related works on API invocations mostly rely on variations of the aforementioned standard approaches. Some works use humans [42] or LLMs [39] as judges, and others execute the API invocation and verify the state of the environment [30, 41]. We are not aware of an evaluation method able to provide results as fine-grained as ours.
@@
LLMs display a remarkable capacity to memorize information from their training data—which can be beneficial when it comes to generating factual responses, but also a problem when it comes to leaking private information from the training data. Studies have shown that LLMs memorize, among other things, method definitions and method calls [43], but also that they struggle to suggest the correct APIs [11]. These findings are in line with our observation that LLMs seem to know many endpoint URLs and argument names, but are often unable to assemble them to a specification-compliant API invocation that solves the task in question.
If LLMs are unable to recall memorized facts, they tend to hallucinate information that is not grounded in reality. This is a common error of LLMs, and studies have shown that it also exists in code generation, where, among other things, models may hallucinate function names and arguments [44, 6, 7]. Our study confirms that these findings also hold for web API endpoints and their arguments, where a high amount of hallucinations can be observed.
@@
Correct usage of APIs is challenging also for humans, as indicated by several studies on API usability issues [3, 4] and real-world API misuses [45, 46, 47]. Several (static analysis) methods have been proposed for detecting (and mitigating) API misuses, e.g., [45, 48, 47]. They are, however, primarily designed to aid humans, not LLMs. Using static analyses for evaluation purposes is generally conceivable, but they can only detect whether an API invocation is legal or illegal, not whether it also solves the given task.
V Conclusion and Future Work
@@
In this paper, we examined the challenges faced by large language models when generating reliable web API invocation code. We introduced WAPIIBench, a novel dataset and evaluation pipeline to systematically assess LLM performance in this domain. Our experiments with open-source LLMs revealed significant limitations: Only 30% (full completion) and 40% (argument completion) of the produced samples were completely right. Endpoints and arguments are often hallucinated due to insufficient recall of knowledge on API usage.
@@
Overall, our work underscores the need for integrating quality assurance techniques into AI-driven code generation workflows to ensure correctness and reliability. Our benchmark lays the foundation for future work to investigate the impact of different approaches—such as retrieval-augmented generation [49], fine-tuning, reasoning, feedback loops, and constrained decoding [50, 51]—on the correctness and specification-compliance of generated web API invocation code. While our evaluation pipeline is specialized on JavaScript and Axios, our dataset and methodology is generic and can be transferred to other languages and libraries.
Acknowledgments